# de Valparaíso CCTVal, Casilla 110-V, Valparaíso, Chile

### v1

@@ 1 Introduction
A number of variations of the scotogenic model have also been discussed. Examples are the “scoto-singlet” model (adding a $S_{1,1,0}$)  [^2] We will use $S$ ($F$) for a scalar (fermion) BSM field, with the subscripts denote the properties of the field under the SM gauge group $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$. Beniwal:2020hjc or the type-I/III variant (adding a fermionic triplet, $F_{1,3,0}$) Hirsch:2013ola . A version with two inert doublet scalars (and only one $N$) was presented in Hagedorn:2018spx . A generalization of the scotogenic model to an arbitrary number of generations (both scalars and fermions) was discussed in Escribano:2020iqq . Adding a $Z^{\prime}$ to the model allows to motivate the $Z_{2}$ (which is put by hand in the orignal model) and affects drastically the relic density of the $N$ in the scotogenic model, as discussed in Chun:2023vbh . A general list of UV extensions that could explain the $Z_{2}$ of the scotogenic model was given in Portillo-Sanchez:2023kbz . Spontaneous breaking of lepton number in a scotogenic setup was considered in DeRomeri:2022cem . All these variations, however, keep $\eta$ and $N_{i}$ as the basic ingredient of the models.
@@ 2.1 Model lagrangian and mass matrices
$h_{F}=\frac{c}{2{\cal F}}\sqrt{{\hat{M}}_{R}^{-1}}h_{\bar{F}}^{-1}U_{\nu}^{*}\hat{m_{\nu}}U_{\nu}^{\dagger},$ (10)
where $[M_{R}]=[{\rm GeV}]$, $[m_{\nu}]=[{\rm eV}]$ and $c=10^{-9}[{\rm GeV}]^{-1/2}$, while the Yukawa couplings $h_{F}$ and $h_{\bar{F}}$ are adimensional. Here, $\hat{m_{\nu}}$ are the light neutrino mass eigenvalues ($m_{\nu_{i}}={m_{\nu_{1}}}$, $\sqrt{\Delta(m_{sol}^{2})+m_{\nu_{1}}^{2}}$, $\sqrt{\Delta(m_{atm}^{2})+m_{\nu_{1}}^{2}}$), with $\Delta(m_{sol}^{2})$ and $\Delta(m_{atm}^{2})$ the solar and atmospheric mass squared splittings and $U_{\nu}$ is the neutrino mixing matrix, measured in oscillation experiments.
@@ 3 Dark matter phenomenology
When the mass splitting between the singlet and triplet scalars is not significantly large, however, co-annihilation channels involving $t^{++}$, $t^{+}$, $t^{0}$ come into play and dominate $\langle\sigma v\rangle$. The importance of co-annihilation processes becomes substantial when the relative mass splittings of the triplet scalars with $S_{0}$ are much less than $1$ Griest:1990kh . In this case, $\lambda_{5}$ and $\Delta m$ become the key parameters in determining the DM relic abundance, allowing the correct $\Omega h^{2}$ to be obtained in a broader range of $S_{0}$ masses, thanks to the more efficient annihilation.
@@ Acknowledgements
G.C. thanks Gabriel Torrealba for his help with python and Andre Lessa for useful comments. M.H. is supported by the Spanish grants PID2020-113775GB-I00 (AEI/10.13039/ 501100011033), CIPROM/2021/054 (Generalitat Valenciana) and the MultiDark network, RED2022-134411-T. G.C. acknowledges support from ANID FONDECYT grant No. 11220237. J.C.H. acknowledges support from ANID FONDECYT grant No.1241685. T.B.M acknowledges support from ANID FONDECYT grant No. 3220454. G.C., J.C.H. and T.B.M. also acknowledge support from ANID – Millennium Science Initiative Program ICN2019_044. C.A. is supported by ANID-Chile FONDECYT grant No. 1231248 and ANID-Chile PIA/APOYO AFB230003.

### v2

@@ 1 Introduction
A number of variations of the scotogenic model have also been discussed. Examples are the “scoto-singlet” model (adding a $S_{1,1,0}$)  [^2] We will use $S$ ($F$) for a scalar (fermion) BSM field, with the subscripts denote the properties of the field under the SM gauge group $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$. Beniwal:2020hjc or the type-I/III variant (adding a fermionic triplet, $F_{1,3,0}$) Hirsch:2013ola . A version with two inert doublet scalars (and only one $N$) was presented in Hagedorn:2018spx . A generalization of the scotogenic model to an arbitrary number of generations (both scalars and fermions) was discussed in Escribano:2020iqq . Adding a $Z^{\prime}$ to the model allows to motivate the $Z_{2}$ (which is put by hand in the orignal model) and affects drastically the relic density of the $N$ in the scotogenic model, as discussed in Kanemura:2011vm . A general list of UV extensions that could explain the $Z_{2}$ of the scotogenic model was given in Portillo-Sanchez:2023kbz . Spontaneous breaking of lepton number in a scotogenic setup was considered in DeRomeri:2022cem . All these variations keep $\eta$ and $N_{i}$ as the basic ingredient of the models. However, there is also Lineros:2020eit , which uses two scalar triplets instead of $\eta$ and Farzan:2010mr with a scalar triplet and a scalar singlet.
@@ 2.1 Model lagrangian and mass matrices
$h_{F}=\frac{1}{2{\cal F}}{\hat{M}}_{R}^{-1}(h_{\bar{F}}^{T})^{-1}U_{\nu}^{*}\hat{m_{\nu}}U_{\nu}^{\dagger},$ (10)
where $[M_{R}]=[{\rm GeV}]$, $[m_{\nu}]=[{\rm eV}]$, while the Yukawa couplings $h_{F}$ and $h_{\bar{F}}$ are adimensional. Here, $\hat{m_{\nu}}$ are the light neutrino mass eigenvalues ($m_{\nu_{i}}={m_{\nu_{1}}}$, $\sqrt{\Delta(m_{sol}^{2})+m_{\nu_{1}}^{2}}$, $\sqrt{\Delta(m_{atm}^{2})+m_{\nu_{1}}^{2}}$), with $\Delta(m_{sol}^{2})$ and $\Delta(m_{atm}^{2})$ the solar and atmospheric mass squared splittings and $U_{\nu}$ is the neutrino mixing matrix, measured in oscillation experiments.
@@ 3 Dark matter phenomenology
When the mass splitting between the singlet and triplet scalars is not significantly large, however, co-annihilation channels involving $t^{++}$, $t^{+}$, $t^{0}$ come into play and dominate $\langle\sigma v\rangle$. The importance of co-annihilation processes becomes substantial when the relative mass splittings of the triplet scalars with $S_{0}$ are much less than $1$ Griest:1990kh . In this case, $\lambda_{5}$ and $\Delta m$ become the key parameters in determining the DM relic abundance. This is demonstrated in fig. (4) for some specific example. Co-annihilation allows the correct $\Omega h^{2}$ to be obtained in a broader range of $S_{0}$ masses.
Figure 4: The plot shows $\Omega h^{2}$ as function of $\Delta m$ for four different choices of $\lambda_{5}$ for one fixed, but arbitrary choice of $m_{S_{0}}=1$ TeV. Smaller values of $\lambda_{5}$ lead to smaller mixing between the neutral singlet and triplet states and, thus to a larger $\Omega h^{2}$. However, even for very small $\lambda_{5}$, the correct $\Omega h^{2}$ can be achieved if $\Delta m$ is small. The plot therefore demonstrates the importance of coannihilation.
@@ 5 Summary and discussion
This paper is mostly concerned with the determination of the parameter space, where the $t^{++}$ predicted by our model could be discovered at the LHC. Before closing, we would like to discuss briefly the question, whether such a discovery could actually be used to demonstrate that the neutrino mass generation mechanism has been identified experimentally.
All Majorana neutrino mass models violate lepton number. Thus, it seems straightforward: Demonstrate that the particle discovered has both, a lepton number conserving (LNC) and a lepton number violating (LNV) decay mode. However, given the finite statistics expected even for ${\cal L}=3/$ab, this requires that the hadronic and leptonic final states of the $t^{++}$ decay have very similar branching ratios, otherwise only one of the two will be observable. As fig. (3) shows, parameter points with $\Gamma_{ll}=\Gamma_{had}$ exist. However, it is also obvious that the ratio of these two widths can differ widely. Thus, the discovery of lepton number violation is not guaranteed, even if the $t^{++}$ is found at the LHC.
Even if both decay modes were measured at the LHC, there is the problem that all final states contain a DM particle, i.e. experimentally missing energy. LNV can not be easily established in events with missing energy. To demonstrate that the missing energy is not due to emission of neutrinos (which would make the leptonic width actually a LNC mode) in the event is far from trivial too: One would need to establish either that the escaping particle is massive (SM neutrinos are “massless” for the LHC) or demonstrate (for example, by studying the energy distributions of the charged leptons) that only a single particle has escaped detection (and not two, as would be necessary to compensate the lepton number of the two charged leptons). In summary, it seems possible, in principle, to show experimentally that the LHC has found LNV and thus traces of the origin of neutrino mass. However, in practice to make all necessary measurements requires that the model is realized in nature within some particularly favourable part of parameter space.
@@ Acknowledgements
G.C. thanks Gabriel Torrealba for his help with python and Andre Lessa for useful comments. M.H. is supported by the Spanish grants PID2020-113775GB-I00 (AEI/10.13039/ 501100011033), CIPROM/2021/054 (Generalitat Valenciana) and the MultiDark network, RED2022-134411-T. G.C. acknowledges support from ANID FONDECYT grant No. 11220237. J.C.H. acknowledges support from ANID FONDECYT grant No.1241685. T.B.M acknowledges support from ANID FONDECYT grant No. 3220454. G.C., J.C.H. and T.B.M. also acknowledge support from ANID - Millennium Science Initiative Program ICN2019_044. C.A. is supported by ANID-Chile FONDECYT grant No. 1231248 and ANID-Chile PIA/APOYO AFB230003.